OpenStudy (anonymous):

Use the ratio test to find the radius of convergence and interval of convergence of the power series.

OpenStudy (anonymous):

\[x-(x^2/4)+(x^3/9)-(x^4/16)+(x^5/25)-...\]

OpenStudy (anonymous):

this the given that you write? x−(x2/4)+(x3/9)−(x4/16)+(x5/25)−...

OpenStudy (anonymous):

yes this is the series

OpenStudy (anonymous):

how do i find the radius of convergence of this power series ?

OpenStudy (anonymous):

you are not giving the answer this different series....

OpenStudy (anonymous):

what?

terenzreignz (terenzreignz):

Do you know how to use the Ratio Test?

OpenStudy (anonymous):

math give a link that not helpful for you he giving you answer for different series

OpenStudy (anonymous):

@nader1 im helping with similar problem

OpenStudy (anonymous):

yes i know how to use the ratio test

OpenStudy (anonymous):

how do i find the sigma notation of this series?

terenzreignz (terenzreignz):

Say you have a series \[\huge \sum_{n=1}^\infty a_n\]Then if you consider the limit \[\huge k=\lim_{n \rightarrow \infty}\left|\frac{a_{n+1}}{a_n}\right|\] Then if k > 0, then it's divergent k < 0, then it's absolutely convergent k = 1, no conclusion may be drawn

terenzreignz (terenzreignz):

Now as for a sigma notation, look for a pattern... all power series take the form \[\huge \sum_{n=1}^\infty a_nx^n\]right?

terenzreignz (terenzreignz):

So, let's examine your series... \[\Large x -\frac{x^2}{4}+\frac{x^3}{9}-\frac{x^4}{16}+\frac{x^5}{25}-...\] see a pattern?

OpenStudy (anonymous):

x^n/n^2?

terenzreignz (terenzreignz):

Not quite :) As you can see, the signs alternate.

OpenStudy (anonymous):

\[(-1)^{n-1} \frac{ x^n }{ n^2 }\]

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

Yeah, that's it :) \[\huge \sum_{n=1}^\infty (-1)^{n-1}\frac{x^n}{n^2}\]

terenzreignz (terenzreignz):

Not that the -1 really matters, you're going to take the absolute value anyway, when you use the ratio test. Please do that now ^.^

OpenStudy (anonymous):

i do the ratio test now?

terenzreignz (terenzreignz):

Yes.

OpenStudy (anonymous):

ugh is a long one

OpenStudy (anonymous):

is there a trick to do it faster

terenzreignz (terenzreignz):

Not really :) After all, you're getting absolute value, means that that \[(-1)^{n+1}\]won't mean squat when you're using the Ratio test.

terenzreignz (terenzreignz):

So, using the ratio test, we need this limit \[\huge k = \lim_{n\rightarrow\infty}\left|\frac{\frac{x^{n+1}}{(n+1)^2}}{\frac{x^n}{n^2}}\right|\]

OpenStudy (anonymous):

can you reduce it to \[\frac{ x*(n+1)^2 }{ n^2 }\]

terenzreignz (terenzreignz):

That you can :)

terenzreignz (terenzreignz):

Don't forget absolute value, ok?

OpenStudy (anonymous):

so limit is x?

terenzreignz (terenzreignz):

Almost, actually, it's |x|

OpenStudy (anonymous):

oh yea

terenzreignz (terenzreignz):

So, can you do it from here?

OpenStudy (anonymous):

how do i find the radius from here?

terenzreignz (terenzreignz):

The limit k in your application of the Ratio Test was |x| right? Well, what must that limit be, if your series is to be convergent?

OpenStudy (anonymous):

1?

terenzreignz (terenzreignz):

That's right, or less :) So, \[\large |x|\le1\]There's your radius of convergence

terenzreignz (terenzreignz):

Whoops, sorry \[\large |x| < 1\]

OpenStudy (anonymous):

thank you

terenzreignz (terenzreignz):

No problem :)